# Universal Anomaly Detection at the LHC: Transforming Optimal Classifiers and the DDD Method

### v2

@@ 1 Introduction
The Standard Model (SM) of particle physics cannot (unfortunately) describe the complete world. It is scientifically accepted that phenomena such as dark matter, baryon asymmetry, neutrino masses, and the description of gravity in regimes where quantum effects are significant require extensions beyond the SM and/or General Relativity. This extension may give rise to new particles and/or effects that are visible in the data from particle physics experiments.
@@
Assuming a signal and supervised event classifiers: The search for these effects is like looking for a needle in a haystack. The effects of new particles generally have a small cross-section and have to be filtered out of an enormous number of background events. Classification algorithms have been used for this purpose since the 1980s or 1990s, see e.g. Peterson:1993nk ; Stimpfl-Abele:1993xgr ; Ametller:1996ri ; Denby:1987rk and since 2014 Baldi:2014kfa , more and more deep learning algorithms have been used. Classification algorithms such as neural networks learn with labelled data (i.e. supervised) to distinguish signals and background events given physical features (such as the 4-vectors and object types) based on training data by adjusting weights to minimize a so-called loss function. Typically, the loss function is a binary cross entropy, that measures the difference of the true class probability distribution to the predicted class probability distribution. Minimizing this penalty function then leads to the classifier learning the ratio of probability distributions in the feature space between signal and background. In recent years, deep learning architectures have been shown to provide better classifiers than the boosted decision trees widely used in high-energy physics. More recently, fully connected deep learning architectures have been overtaken by graph networks and even attention-based transformer architectures, inspired by their success in language models.
@@
Not assuming a signal and unsupervised anomaly detection: However, the question arises “what is the signal ?”. As mentioned above, the exact signal of physics beyond the SM (BSM) is not known, so a signal-by-signal search may not be efficient. One search method that has gained more and more attention in recent decades is the systematic search for anomalies, i.e., the model-independent search for events that are not SM-like. One method is to test numerous signal regions using automated procedures D0:2000vuh ; Abbott:2000gx ; Abbott:2001ke ; Abazov:2011ma ; H1:2004rlm ; H1:2008aak ; CDF:2007iou ; CDF:2008voc ; Aaboud:2018ufy ; Sirunyan:2020jwk . Other approaches use unsupervised (or weakly supervised) classifiers. Machine Learning (ML) algorithms are trained without a signal model to estimate the distribution of SM events. Unsupervised refers to no signal data being used to train the ML model. Many approaches use AutoEncoders, mapping data features $x$ to a code $z$ and back to $x^{\prime}$, assuming poor reconstruction of anomalies. In extreme unsupervised cases, only the expected density of the Standard Model is needed Caron:2021wmq , derived from simulated events, direct data (assuming a small signal), or auxiliary measurements. Unsupervised Deep Learning techniques to search for new physics are discussed in  Wozniak:2020cry ; Knapp:2020dde ; Dery:2017fap ; Cohen:2017exh ; Metodiev:2017vrx ; DAgnolo:2018cun ; Heimel:2018mkt ; Farina:2018fyg ; Komiske:2018oaa ; Hajer:2018kqm ; Cerri:2018anq ; Collins:2018epr ; Otten:2019hhl ; Blance:2019ibf ; Collins:2019jip ; Sirunyan:2019jud ; vanBeekveld:2020txa ; Aad:2020cws ; Nachman:2020lpy ; Kasieczka:2021xcg ; Caron:2021wmq ; Aad_2024 ; CMS-PAS-EXO-22-026 ; Harris:2881089 ; Kosters:2022amb . Comparisons of unsupervised approaches are in the Dark Machines challenge Aarrestad_2022 using datasets to compare supervised strategies with unsupervised approaches. Another comparison of model-independent approaches using different black boxes based on data and expectation density comparisons is the LHC Olympics Kasieczka:2021xcg .
@@
The effectiveness of the DDD strategy is compared against leading models from the Dark Machines competition, in particular the DeepSVDD model and a novel model known as Deep Robust One-Class Classification (DROCC), as outlined in Ref goyal2020drocc . The models are evaluated on five different datasets: one is from the search for 4-top production at the LHC ATLAS:2023ajo ; CMS:2023ftu , the other four from the Dark Machines Anomaly Score (DarkMachines) challenge.
@@ 3.1 Multi-Layer Perceptron
In order to interpret the success of a complex architecture performing an anomaly detection task, it is crucial to compare with simpler architectures playing the role of benchmarks goodfellow2016deep . With this goal, a deep MLP implementation have been developed, as illustrated in Table 2.
@@ 3.2 Particle Transformer
$\text{Attention}(Q,K,T)=\text{SoftMax}\left(QK^{T}/\sqrt{d}+U\right)V$ (1)
@@ 4.1 DeepSVDD: Deep Support Vector Data Description
There are infinite possibilities for the choice of a centre in the output space, but the option taken in this study is proposed in Ref. deepSVDD:pmlr-v80-ruff18a . The center is computed as the mean of the network representations that results from an initial forward pass on the training data. This paper also proposes that biases terms need to be removed from the network to avoid trivial solutions. In addition to the previous considerations, not making use of dropout has been shown to be a convenient approach, improving significantly the stability of the training. DeepSVDD methods can typically address this problem with different sets of weights, representing different minima in the weight space. However, some of these minima may lead to ‘sphere collapse’, where the model predicts the output independently of the input. To enhance the stability of this method, an ensemble approach, as proposed in Ref. Caron:2021wmq , is used. This involves computing the mean of the scores from multiple models. Separate models are trained with 2, 4, 8, and 16 output dimensions, respectively, and these models are then combined to form the ensemble.
@@ 4.2 DROCC: Deep Robust One-Class Classification
DROCC goyal2020drocc is a one-class classification method that assumes data from the target class lies on a well-sampled, locally linear, low-dimensional manifold. This assumption enables DROCC to generate synthetic outliers during training to robustly learn a boundary around the target class data. The main idea is to train a classifier that not only accurately classifies the given ‘normal’ data but also is robust enough to classify unseen and potentially very different ‘anomalous’ data as being out-of-class.
This is done thought a loss function, which includes a component that actively searches for adversarial examples near the ‘normal’ data points. These adversarial examples are synthetic data points that are slightly perturbed versions of the normal data and are hypothesized to lie close to the decision boundary between normal and anomalous classes. The objective is to ensure that the model can correctly classify these challenging examples as anomalies, thereby enhancing its robustness to real anomalous data it may encounter. In this way, avoids the phenomenon of ‘sphere collapse’.
In the proposed approach, the model training procedure uses the test data with signals to optimize the training process by maximizing the Area Under the Curve (AUC). This inclusion of test data makes the method not entirely unsupervised (it can be considered as a form of weak supervision), as it stabilizes the otherwise unstable training process. The AUC provides a more stable metric that reflects the model’s ability to distinguish between normal and anomalous data points.
Table 4 lists the hyperparameter settings used for training the DROCC model:
Parameter Value Description
$r$ 4 Radius for synthetic negative examples
$\gamma$ 1 Scaling factor for radius
$\mu$ 0.5 Weight for synthetic negative examples in the loss function
Learning Rate (GA) 0.5 Learning rate for gradient ascent
Learning Rate (GD) 0.001 Learning rate for gradient descent
Table 4: Hyperparameter settings for the DROCC model.
@@ 4.3 DDD: Discriminatory detection of distortions
Algorithm Steps
@@
1:Data Preparation:
2:Load the background dataset $D_{\text{background}}$.
3:Split $D_{\text{background}}$ into two sets: $D_{\text{original}}$ (e.g., even events) and $D_{\text{mod\_source}}$ (e.g., odd events).
4:Data Modification:
5:for each event in $D_{\text{mod\_source}}$ do
6:     Apply the modification function $f_{\text{mod}}(\theta)$, where $\theta\sim N(1,\text{std})$, with std ranging from 0.00 to 0.10, to generate the modified dataset $D_{\text{modified}}$.
7:     for each object in the event do
8:         With probability $p$, randomly decide to add a new object.
9:         With the same probability $p$, randomly decide to remove an existing object.
10:     end for
11:     Adjust $\theta$ within the specified standard deviation range for subsequent modifications.
12:end for
13:Discriminator Model Definition:
14:Define the discriminator model $M_{\text{discriminator}}$.
15:Training:
16:Combine $D_{\text{original}}$ and $D_{\text{modified}}$ into a combined dataset $D_{\text{combined}}$.
17:Label $D_{\text{original}}$ as ‘0’ and $D_{\text{modified}}$ as ‘1’.
18:Train $M_{\text{discriminator}}$ on $D_{\text{combined}}$.
19:Evaluation:
20:Calculate the AUC score using predictions on $D_{\text{original}}$ and $D_{\text{modified}}$.
21:Hyperparameter Adjustment:
22:if AUC score deviates significantly from the typical signal value (0.85 as mentioned before) then
23:     Adjust $\theta$ accordingly.
24:     Repeat the process from Step 2 until the desired AUC score is achieved.
25:end if
@@ 5 Results
Figure 3: This figure shows the performance of the eight models on the four channels of the DarkMachines dataset: Channel 1 (first row, left), Channel 2a (first row, right), Channel 2b (second row, left) and Channel 3 (second row, right). For each channel, different signals are evaluated. The performance metrics shown are the AUC and the signal efficiency ($\epsilon_{S}$) assuming a background efficiency ($\epsilon_{B}$) of 1%. A color code is used to distinguish among the three training techniques: DeepSVDD (blue), DROCC (green) and DDD (red). Different shapes are used for each architecture: a circle for the MLP, a square for the ParT with no interactions, and a triangle for the ParT with SM couplings.
@@ 5.1 Performance for the DarkMachines challenge channels
The observed trends are remarkably similar. The results for the individual channels are summarized as follows (numerical details of the measured variables used in this work can be found in the Appendix):
@@ 5.2 Performance for the 4-top channel
Figure 4: Performance of the eight models on the 4-top channel. The performance metrics shown are the AUC and the signal efficiency ($\epsilon_{S}$) assuming a background efficiency ($\epsilon_{B}$) of 1%. A color code is used to distinguish among the three training techniques: DeepSVDD (blue), DROCC (green), and DDD (red). Different shapes are used for each architecture: a circle for the MLP, a square for the ParT with no interactions, and a triangle for the ParT with SM couplings.
@@
From Figure 4 it is observed that the DDD method with ParT architecture (red triangle) achieves the highest AUC value (0.754) and a significant signal efficiency ($\epsilon_{S}$) of 0.033 at $\epsilon_{B}$ = 0.01. This highlights the strength of the ParT architecture in handling complex correlations through its attention mechanism, which effectively avoids bias from padding.
The DROCC technique also shows robust performance, particularly with the ParT architecture, achieving an AUC of 0.75 and $\epsilon_{S}$ of 0.084 at $\epsilon_{B}$ = 0.01. The MLP with DROCC also performs well, with an AUC of 0.71, but slightly lower signal efficiency compared to the ParT architecture.
The DeepSVDD method demonstrates the lowest AUC values among the techniques but still provides competitive results, especially for the MLP architecture with an AUC of 0.606 and a relatively high $\epsilon_{S}$ of 0.124 at $\epsilon_{B}$ = 0.01. The ParT architecture with DeepSVDD shows lower performance, with an AUC of 0.518 and $\epsilon_{S}$ of 0.007 at $\epsilon_{B}$ = 0.01, indicating that the simpler MLP architecture might be more suited for certain anomaly detection tasks when using DeepSVDD.
@@ 5.3 Comparison to the DarkMachines challenge algorithms
Figure 5 provides a comparative analysis of the TI values for different models, including the results of different models in the DarkMachines challenge and the mixture of theories from Ref. Caron_2023 .

### v3

@@ 1 Introduction
The Standard Model (SM) of particle physics cannot describe the complete world. It is scientifically accepted that phenomena such as dark matter, baryon asymmetry, neutrino masses, and the description of gravity require extensions beyond the SM and/or General Relativity. This extension may give rise to new particles and/or effects that are visible in the data from particle physics experiments.
@@
Assuming a signal and supervised event classifiers: The search for these effects is like looking for a needle in a haystack. The production of new particles generally have a small cross-section and have to be filtered out of an enormous number of background events. Classification algorithms have been used for this purpose since the 1980s or 1990s, see e.g. Peterson:1993nk ; Stimpfl-Abele:1993xgr ; Ametller:1996ri ; Denby:1987rk and since 2014 Baldi:2014kfa , more and more deep learning algorithms have been used. Classification algorithms such as neural networks learn with labelled data (i.e. supervised) to distinguish signals and background events given physical features (such as the 4-vectors and object types) based on training data by adjusting weights to minimize a so-called loss function. Typically, the loss function is a binary cross entropy, that measures the difference of the true class probability distribution to the predicted class probability distribution. Minimizing this penalty function then leads to the classifier learning the ratio of probability distributions in the feature space between signal and background. In recent years, deep learning architectures have been shown to provide better classifiers than the boosted decision trees widely used in high-energy physics. More recently, fully connected deep learning architectures have been overtaken by graph networks Shlomi:2020gdn and even attention-based transformer architectures qu2024particle , inspired by their success in language models DBLP:journals/corr/VaswaniSPUJGKP17 .
@@
Not assuming a signal and unsupervised anomaly detection: However, the question arises “which signal is considered?”. As mentioned above, the exact signal of physics beyond the SM (BSM) is not known, so a signal-by-signal search may not be efficient. One search method that has gained more and more attention in recent decades is the systematic search for anomalies, i.e., the model-independent search for events that are not SM-like. One method is to test numerous signal regions using automated procedures D0:2000vuh ; Abbott:2000gx ; Abbott:2001ke ; Abazov:2011ma ; H1:2004rlm ; H1:2008aak ; CDF:2007iou ; CDF:2008voc ; Aaboud:2018ufy ; Sirunyan:2020jwk . Other approaches use unsupervised (or weakly supervised) classifiers. Machine Learning (ML) algorithms are trained without a signal model to estimate the distribution of SM events. Unsupervised refers to no signal data being used to train the ML model. Many approaches use AutoEncoders, mapping data features $x$ to a code $z$ and back to $x^{\prime}$, assuming poor reconstruction of anomalies. In extreme unsupervised cases, only the expected density of the Standard Model is needed Caron:2021wmq , derived from simulated events, direct data (assuming a small signal), or auxiliary measurements. Unsupervised Deep Learning techniques to search for new physics are discussed in Wozniak:2020cry ; Knapp:2020dde ; Dery:2017fap ; Cohen:2017exh ; Metodiev:2017vrx ; DAgnolo:2018cun ; Heimel:2018mkt ; Farina:2018fyg ; Komiske:2018oaa ; Hajer:2018kqm ; Cerri:2018anq ; Collins:2018epr ; Otten:2019hhl ; Blance:2019ibf ; Collins:2019jip ; Sirunyan:2019jud ; vanBeekveld:2020txa ; Aad:2020cws ; Nachman:2020lpy ; Kasieczka:2021xcg ; Caron:2021wmq ; Aad_2024 ; CMS-PAS-EXO-22-026 ; Harris:2881089 ; Kosters:2022amb . Comparisons of unsupervised approaches are in the Dark Machines challenge Aarrestad_2022 using datasets to compare supervised strategies with unsupervised approaches. Another comparison of model-independent approaches using different black boxes based on data and expectation density comparisons is the LHC Olympics Kasieczka:2021xcg .
@@
The effectiveness of the DDD strategy is compared against leading models from the Dark Machines competition, in particular the DeepSVDD model deepSVDD:pmlr-v80-ruff18a and a novel model known as Deep Robust One-Class Classification (DROCC), as outlined in Ref goyal2020drocc . The models are evaluated on five different datasets: one is from the search for 4-top production at the LHC ATLAS:2023ajo ; CMS:2023ftu , the other four from the Dark Machines Anomaly Score (DarkMachines) challenge.
@@ 3.1 Multi-Layer Perceptron
In order to interpret the success of a complex architecture performing an anomaly detection task, it is crucial to compare with simpler architectures playing the role of benchmarks goodfellow2016deep . With this goal, a deep MLP implementation has been developed, as illustrated in Table 2.
@@ 3.2 Particle Transformer
$\text{Attention}(Q,K,V)=\text{SoftMax}\left(QK^{T}/\sqrt{d}+U\right)V$ (1)
@@ 4.1 DeepSVDD: Deep Support Vector Data Description
There are infinite possibilities for the choice of a centre in the output space, but the option taken in this study is proposed in Ref. deepSVDD:pmlr-v80-ruff18a . The center is computed as the mean of the network representations that results from an initial forward pass on the training data. Removing all bias terms from the network is also needed to avoid trivial solutions. In addition to these considerations, not making use of dropout has been shown to be a convenient approach, improving significantly the stability of the training.
DeepSVDD methods can typically address this problem with different sets of weights, representing different minima in the weight space. However, some of these minima may lead to ‘sphere collapse’, where the model predicts the output independently of the input. To enhance the stability of this method, an ensemble approach, as proposed in Ref. Caron:2021wmq , is used. This involves computing the mean of the scores from multiple models. Separate models are trained with 2, 4, 8, and 16 output dimensions, respectively, and these models are then combined to form the ensemble.
Any classifier can be transformed into an anomaly detector taking into account the considerations described above. The procedure is summarized in Algorithm 1.
Step 1: Remove all bias terms from the network;
Step 2: Avoid the use of dropout layers;
Step 3: Add an output layer with a certain number of nodes, which will define the dimension of the output hyperspace;
Step 4: Re-define the loss function as the distance to a target center in the hyperspace, which needs to be minimized during the training on the background events;
Step 5: Define the center of the hypersphere as the mean of the output from an initial forward pass on the training data;
Step 6: To improve stability, train separate models with different output dimensions (e.g., 2, 4, 8, and 16 nodes in this case). Take the mean scores for each event.
Algorithm 1 Transforming a Classifier into an Anomaly Detector using the DeepSVDD technique
@@ 4.2 DROCC: Deep Robust One-Class Classification
DROCC goyal2020drocc  [^1] To the best of our knowledge, this is the first application of the DROCC method in the context of LHC physics. is a one-class classification method that assumes data from the target class lies on a well-sampled, locally linear, low-dimensional manifold. This assumption enables DROCC to generate synthetic outliers during training to robustly learn a boundary around the target class data. The main idea is to train a classifier that not only accurately classifies the given ‘normal’ data but also is robust enough to classify unseen and potentially very different ‘anomalous’ data as being out-of-class.
This is done through the loss function, which includes a component that actively searches for adversarial examples near the ‘normal’ data points. These adversarial examples are synthetic data points that are slightly perturbed versions of the normal data and are hypothesized to lie close to the decision boundary between normal and anomalous classes. The objective is to ensure that the model can correctly classify these challenging examples as anomalies, thereby enhancing its robustness to real anomalous data it may encounter. In this way, avoids the phenomenon of ‘sphere collapse’.
In the proposed approach, training is conducted entirely with background (target class) events to compute the cross-entropy loss and the adversarial loss. Signals are introduced only during the validation phase, where they are used to compute evaluation metrics, such as the Area Under the Curve (AUC). These signals, however, are not part of the training process and are distinct from those used in the final evaluation after training. The signals used for validation are taken from Ref. Caron_2023 , whereas those employed during the final test phase are from Re. Aarrestad_2022 . This weak supervision approach is motivated by the instability observed in the AUC during training. Monitoring the AUC after each epoch reveals significant fluctuations, and stopping training arbitrarily can result in poor performance at the final evaluation. By incorporating signals in the evaluation phase, we stabilize the process and guide hyperparameter optimization, ensuring that the model converges to a more robust solution. This allows the DROCC method to remain a one-class classification technique during training, while still benefiting from signals in the validation phase to improve AUC stability and overall model performance.
Algorithm 2 provides a step-by-step outline of the DROCC method.
Training Phase:
for $\text{epoch}=1$ to Max_epochs do
       for batch in $X_{\text{background}}$ do
             Generate adversarial examples $X_{\text{adv}}$ for $X_{\text{bkg}}$:
              Apply small random perturbations to $X_{\text{bkg}}$;
              Project perturbed samples onto a Hypersphere constraint;
             Compute loss:
              $L_{\text{original}}=\text{CrossEntropy}(f_{\text{model}}(X_{\text{bkg}}),\text{BackgroundLabel})$;
              $L_{\text{adv}}=\text{CrossEntropy}(f_{\text{model}}(X_{\text{adv}}),\text{AnomalyLabel})$;
              $\text{TotalLoss}=L_{\text{original}}+L_{\text{adv}}$;
             Update model parameters $\theta$ using gradient descent to minimize the TotalLoss;
      Validation Phase:
       Evaluate $f_{\text{model}}$ on validation set $(X_{\text{val\_bkg}},X_{\text{val\_sig}})$;
       Compute the AUC to monitor overfitting and guide early stopping;
Testing Phase:
Evaluate $f_{\text{model}}$ on test set $(X_{\text{test\_bkg}},X_{\text{test\_sig}})$;
Compute final metrics.
Algorithm 2 DROCC for Anomaly Detection. For more details, see goyal2020drocc .
@@ 4.3 DDD: Discriminatory detection of distortions
The DDD method is summarized in Algorithm 3.
Step 1: Data Preparation
Load the background dataset $D_{\text{background}}$;
Split $D_{\text{background}}$ into two sets: $D_{\text{original}}$ (e.g., even events) and $D_{\text{mod\_source}}$ (e.g., odd events);
Step 2: Data Modification
for each event in $D_{\text{mod\_source}}$ do
       Apply the modification function $f_{\text{mod}}(\theta)$, where $\theta\sim N(1,\text{std})$, with std ranging from 0.00 to 0.10, to generate the modified dataset $D_{\text{modified}}$;
       for each object in the event do
             With probability $p$, randomly decide to add a new object;
             With the same probability $p$, randomly decide to remove an existing object;
       end for
      Adjust $\theta$ within the specified standard deviation range for subsequent modifications;
end for
Step 3: Discriminator Model Definition
Define the discriminator model $M_{\text{discriminator}}$;
Step 4: Training
Combine $D_{\text{original}}$ and $D_{\text{modified}}$ into a combined dataset $D_{\text{combined}}$;
Label $D_{\text{original}}$ as ‘0’ and $D_{\text{modified}}$ as ‘1’;
Train $M_{\text{discriminator}}$ on $D_{\text{combined}}$;
Step 5: Evaluation
Calculate the AUC score using predictions on $D_{\text{original}}$ and $D_{\text{modified}}$;
Step 6: Hyperparameter Adjustment
if AUC score deviates significantly from the typical signal value (0.85 as mentioned) then
       Adjust $\theta$ accordingly;
       Repeat the process from Step 2 until the desired AUC score is achieved.
end if
@@ 5 Results
Figure 3: This figure shows the performance of the eight models on the four channels of the DarkMachines dataset: Channel 1 (first row, left), Channel 2a (first row, right), Channel 2b (second row, left) and Channel 3 (second row, right). For each channel, different signals are evaluated. The performance metrics shown are the AUC and the signal efficiency ($\epsilon_{S}$) assuming a background efficiency ($\epsilon_{B}$) of 1%. A vertical line is added to indicate a significance improvement equal to 1. A color code is used to distinguish among the three training techniques: DeepSVDD (blue), DROCC (green) and DDD (red). Different shapes are used for each architecture: a circle for the MLP, a square for the ParT with no interactions, and a triangle for the ParT with SM couplings.
@@ 5.1 Performance for the DarkMachines challenge channels
The observed trends are remarkably similar. The results for the individual channels are summarized as follows (numerical details of the measured variables used in this work can be found in Appendix A.1):
@@ 5.2 Performance for the 4-top channel
Figure 4: Performance of the eight models on the 4-top channel. The performance metrics shown are the AUC and the signal efficiency ($\epsilon_{S}$) assuming a background efficiency ($\epsilon_{B}$) of 10%. A vertical line is added to indicate a significance improvement equal to 1. A color code is used to distinguish among the three training techniques: DeepSVDD (blue), DROCC (green), and DDD (red). Different shapes are used for each architecture: a circle for the MLP, a square for the ParT with no interactions, and a triangle for the ParT with SM couplings.
@@
From Figure 4 it is observed that the DDD method with ParT architecture (red triangle) achieves the highest AUC value (0.754) and a significant signal efficiency ($\epsilon_{S}$) of 0.307 at $\epsilon_{B}$ = 0.1. This highlights the strength of the ParT architecture in handling complex correlations through its attention mechanism, which effectively avoids bias from padding.
The DROCC technique also shows robust performance, particularly with the MLP architecture, achieving an AUC of 0.71 and $\epsilon_{S}$ of 0.230 at $\epsilon_{B}$ = 0.1. The performance is slightly lower for the ParT implementations.
The DeepSVDD method demonstrates the lowest AUC values among the techniques but still provides competitive results, especially for the MLP architecture with an AUC of 0.606 and a relatively high $\epsilon_{S}$ of 0.237 at $\epsilon_{B}$ = 0.1. The ParT architecture with DeepSVDD shows lower performance, with an AUC of 0.518 and $\epsilon_{S}$ of 0.103 at $\epsilon_{B}$ = 0.1, indicating that the simpler MLP architecture might be more suited for certain anomaly detection tasks when using DeepSVDD.
@@ 5.3 Comparison to the DarkMachines challenge algorithms
Figure 5 provides a comparative analysis of the TI values for different models, including the results of different models in the DarkMachines challenge and the mixture of theories from Ref. Caron_2023 . The TI is calculated over different background efficiency cuts ($\epsilon_{B}=10^{-2},10^{-3},$ and $10^{-4}$), and the maximum is taken, which typically leads to $TI>1$ values.